OpenStudy (anonymous):

19+5sinx find highest and lowest values

OpenStudy (binary3i):

sinx can take take values between 1 and -1, so the highest is when sin is 1 and lowest is when it is -1 therefore highest is 19+5*(1)=19+5=24 and lowest is 19+5*(-1)=19-5=14